55 pts   The following temperatures were recorded in a rabbit at various times after being inoculated...

  1. 55 pts   The following temperatures were recorded in a rabbit at various times after being inoculated with the rinderpest virus.

Time after injection (hours)       Temperature (0F)

24                                                     102.8

32                                                     104.5

48                                                     106.5

56                                                     107.0

Test the hypothesis that the correlation in significant at the 0.05 level.
Answer #1
( X) ( Y) X^2 Y^2 X*Y
24 102.8 576 10567.8 2467.2
32 104.5 1024 10920.3 3344
48 106.5 2304 11342.3 5112
56 107 3136 11449 5992

calculation procedure for correlation
sum of (x) = ∑x = 160
sum of (y) = ∑y = 420.8
sum of (x^2)= ∑x^2 = 7040
sum of (y^2)= ∑y^2 = 44279.34
sum of (x*y)= ∑x*y = 16915.2
to caluclate value of r( x,y) = covariance ( x,y ) / sd (x) * sd (y)
covariance ( x,y ) = [ ∑x*y - N *(∑x/N) * (∑y/N) ]/n-1
= 16915.2 - [ 4 * (160/4) * (420.8/4) ]/4- 1
= 20.8
and now to calculate r( x,y) = 20.8/ (SQRT(1/4*16915.2-(1/4*160)^2) ) * ( SQRT(1/4*16915.2-(1/4*420.8)^2)
=20.8 / (12.649*1.672)
=0.984
value of correlation is =0.984
coeffcient of determination = r^2 = 0.967
properties of correlation
1. If r = 1 Corrlation is called Perfect Positive Corrlelation
2. If r = -1 Correlation is called Perfect Negative Correlation
3. If r = 0 Correlation is called Zero Correlation
& with above we conclude that correlation ( r ) is = 0.9835> 0 ,perfect positive correlation              

Given that,
value of r =0.984
number (n)=4
null, Ho: ρ =0
alternate, H1: ρ!=0
level of significance, α = 0.05
from standard normal table, two tailed t α/2 =4.303
since our test is two-tailed
reject Ho, if to < -4.303 OR if to > 4.303
we use test statistic (t) = r / sqrt(1-r^2/(n-2))
to=0.984/(sqrt( ( 1-0.984^2 )/(4-2) )
to =7.81
|to | =7.81
critical value
the value of |t α| at los 0.05% is 4.303
we got |to| =7.81 & | t α | =4.303
make decision
hence value of | to | > | t α| and here we reject Ho
ANSWERS
---------------
null, Ho: ρ =0
alternate, H1: ρ!=0
test statistic: 7.81
critical value: -4.303 , 4.303
decision: reject Ho
we have enough evidence to support the claim that the correlation in significant at the 0.05 level.